# Optimization of deterministic photonic graph state generation via local operations

### v1

@@ Acknowledgement
We thank Sophia Economou for helpful discussions. This work was supported in part by MITACS, and the Air Force Office of Scientific Research (AFOSR) under Grant FA9550-22-1-0062. Hoi-Kwong Lo acknowledges research support from NSERC and Innovative Solutions Canada.
@@
A: Exhaustive analysis of small graphs
@@
To showcase the results of the exhaustive analysis, we look into an entanglement class associated with the graph (A) in Fig. S1, i.e., all graph state that are LC equivalent to this graph up to relabeling of the nodes. We have identified multiple intermediate target graphs to utilize in the generation process of any member of this entanglement class, each tailored for optimizing a certain cost parameter. In Fig. S1, examples of a case with the least number of cnots (B), a case with the smallest emitter depth (C), and a case with minimal number of unitary gates (D) are presented. More details for each option are listed in Table S1. We remark that there might be more than one best performing intermediate graphs depending on the cost function used. One can find the maximum improvement for each cost parameter by comparing the best values with the worst ones present in the last column, showing the maximum of the cost parameters over this class when no optimization is done. Overall, any arbitrary cost function can be employed to filter the results of such an exhaustive analysis, e.g, the fidelity of the final state with the target when a proper noise model is employed. The full dataset containing the results from the analysis of all 45 entanglement classes of graph with $n\leq 7$ nodes is available at Ref. [1].
@@
B: Correlation analysis
In a correlation analysis of the local equivalency class of a graph, the goal is to determine whether there exists a significant correlation between a graph metric (a parameter related to the shape of a graph) and a cost parameter of its generating circuit. We examine how changes in the average values of one metric (chosen as the dependent variable) correspond to variations in the other metric (chosen as the independent variable) across the graphs in the orbit or in a sample from it. For instance, in Fig. S2, the independent variable (x-axis) the is number of edges in the graph and the dependent variable (y-axis) is chosen as the average of the circuit gate-count over the graphs sharing the same number of edges in the sample. This diagram is derived for a sample of the LC orbit of a random graph of size $n$=20 with an edge probability of $p$=0.95. The Pearson correlation coefficient [2] for this specific diagram is $94\%$. Repeating the same procedure for 100 random graphs of the same type ($n$=20, $p$=0.95), we find the distribution of the correlation coefficients to be as seen in Fig. S3. This shows an average Pearson correlation of $0.86\pm 0.07$, indicating that for most of the graphs of this type, a strong correlation (over 0.5) holds true.
Figure S3: The distribution of Pearson coefficients for ‘edge count vs gate count’ correlations for 100 trials with different random graphs of size 20 and edge probability 0.95
C: Repeater graph state analysis
Figure S6: Pattern in RGS orbit leading to identification of elementary circuit blocks to realize different shapes. From left to right, in each step a pair of arms from the RGS is folded into the central highly connected component. Circuit block (a) can be used to add a regular arm (a pair of photons: a core node and a leave attached to it) to the graph under preparation, while the circuit block (b) adds a folded arm (a pair of photons: two core nodes connected to all other core nodes except to each other) to the graph. The unitary gate $U$ is $HPHPX$. The number of cnot gates will be reduced as a consequence of using block (b) instead of (a) when increasing the size of the graph.
The correlation analysis results in identifying interesting patterns in the case of repeater graph states. As evident in the Fig. 5 of main text, a subclass of the orbit shows almost perfect inverse linear correlation between the cnot and edge counts. By studying this subclass of the orbit, one can find the sequence of LC operations required so that the graph, and its corresponding circuit structure, moves one step toward the lower-cost region. Such an analysis gives rise to an efficient algorithm to not only generate the correlated subclass of the orbit, but also the full LC equivalency class for any RGS with arbitrary size. The procedure is outlined in detail in Algorithm 1 below.
It is worth noting that determining the LC equivalency of a graph is not a trivial task; in fact, even determining the size of the orbit is a #P-complete problem [4]. In general as one needs to perform a depth or breadth-first search over sequences of local complementation operations on graph nodes, and check for isomorphisms of newly found graphs in each step which is itself an inefficient task. Access to the whole orbit enables one to confirm the results found through correlation analysis and moreover to perform exhaustive optimization with respect to different cost functions.
Figure S6 shows the pattern observed in the shape of the graphs as we move from one point to the next in the correlated part of Fig. 5. By looking into the generating circuit for each case, we were able to find different types of elementary circuit blocks that can be utilized to increase the number of photons in any part of the state. Ultimately, the best (with respect to the use of cnot gates) alternative shape in the orbit and its respective generating circuit can be identified; as depicted in Fig. S7 and Fig. S8.
Algorithm 1 RGS orbit enumeration
Define orbit as an empty list
Add the original RGS $G_{0}$ to the orbit
Define core-list as the list of indices of the core nodes in graph $G_{0}$
Apply LC transformation on one of the core nodes $i$ of $G_{0}$ to get a new graph $G$
Add $G$ to orbit
Remove $i$ from the core-list
while core-list is NOT empty do
     In $G$, apply LC transformation on one of the nodes $j$ in the core-list to get the new graph; set it as $G_{1}$
     Add $G_{1}$ to orbit
     Remove $j$ from core-list
     Add $G_{2}$ to orbit
     if core-list NOT empty then
         In $G_{1}$, apply LC transformation on one of the nodes $k$ in the core-list to get the new graph; set it as $G_{3}$
         Add $G_{3}$ to orbit
         Remove $k$ from core-list
         Set $G\leftarrow G_{3}$
     end if
end while

### v2

@@ Acknowledgement
We thank Sophia Economou for helpful discussions. This work was supported in part by MITACS, and the Air Force Office of Scientific Research (AFOSR) under Grant FA9550-22-1-0062. Benjamin MacLellan acknowledges support from the NSERC Vanier CGS program. Hoi-Kwong Lo acknowledges research support from NSERC, Innovative Solutions Canada, and NRC (National Research Council of Canada) CSTIP Contribution Agreement QSP 081-1.
@@
A: Repeater graph state analysis
The correlation analysis results in identifying interesting patterns in the case of repeater graph states. As evident in the Fig. 3(d) of main text, a subclass of the orbit shows almost perfect inverse linear correlation between the cnot and edge counts. By studying this subclass of the orbit, one can find the sequence of LC operations required so that the graph, and its corresponding circuit structure, moves one step toward the lower-cost region. Such an analysis gives rise to an efficient algorithm to not only generate the correlated subclass of the orbit, but also the full LC equivalency class for any RGS with arbitrary size. The procedure is outlined in detail in Algorithm 1 below.
Figure S1: Seven first graphs in the orbit of a 9-qubit RGS. The numbers on arrows indicate on what node the local complementation should be applied to obtain the next graph. The observed pattern leads to the identification of an algorithm for a full orbit enumeration and being able to find elementary circuit blocks to realize each different shape. On top, from left to right, in each step a pair of arms from the RGS is folded into the central highly connected component. On bottom, from left to right, in each step one of the leaf nodes (blue nodes) gets connected to nodes 0 and 1, and the nodes 0 and 1 alternate between being connected and not connected. The new edges created in each step are represented with bold lines.
It is worth noting that determining the LC equivalency class of a graph is a non-trivial task; in fact, even determining the size of the orbit is a #P-complete problem [1]. In general as one needs to perform a depth or breadth-first search over sequences of local complementation operations on graph nodes, and check for isomorphisms of newly found graphs in each step which is itself an inefficient task. Access to the whole orbit enables one to confirm the results found through correlation analysis and moreover to perform exhaustive optimization with respect to different cost functions.
Figure S1 shows the pattern observed in the orbit and the required local complementation operations to navigate between them. The graphs on top correspond to the first cases in the linearly correlated part of Fig. 3(d). The graphs on bottom represent the uncorrelated part of Fig. 3(d) (the inset). By looking into the generating circuit for each case, one can find different elementary circuit blocks, as shown in the main article, that are to be utilized to add photons to different parts of the graph. Ultimately, the best (with respect to the use of cnot gates) alternative shape in the orbit and its respective generating circuit can be identified by continuing the trends seen in Fig. S1; the optimal case is depicted in Fig. S2 and Fig. S3.
Algorithm 1 RGS orbit enumeration
Define orbit as an empty list
Add the original RGS $G_{0}$ to the orbit
Define core-list as the list of indices of the central nodes (fully connected portion) in graph $G_{0}$
Apply LC transformation (local complementation) on the first node in the core-list of $G_{0}$ to get a new graph $G$. Refer to this node as $i$ hereafter
Add $G$ to orbit
Remove $i$ from the core-list
while core-list is NOT empty do
     In $G$, apply LC transformation on the next first node $j$ in the core-list to get the new graph; set it as $G_{1}$
     Add $G_{1}$ to orbit
     Remove $j$ from core-list
     Add $G_{2}$ to orbit
     if core-list NOT empty then
         In $G_{1}$, apply LC transformation on the next first node $k$ in the core-list to get the new graph; set it as $G_{3}$
         Add $G_{3}$ to orbit
         Remove $k$ from core-list
         Set $G\leftarrow G_{3}$
     end if
end while
B: Graph-to-circuit maps
The optimization outcome is dependant on the specific deterministic graph to circuit map that is used to find a generating circuit for each input graph, however, the framework can take any such graph-to-circuit mapping protocol to work with. The first protocol we implemented (Ref. [27] in the main text) was selected to ensure that the algorithms utilized the minimum number of emitters necessary to generate the entanglement present in a graph state. However, the number of emitters may not always be the primary cost metric and thus other maps can become more suitable options depending on the platform and hardware. For instance, Kaur et al. [2] have shown that considering solely the photon loss as the dominant factor in the failure of an all-photonic repeater protocol, a map that allows for the use of extra emitters (over the theoretical minimum) actually results in a better overall success rate.
Since the optimization scheme proposed here is protocol-agnostic, we analyzed repeater graph states of varying sizes using another specific algorithm (Algorithm 2, as referenced in [2]). In this analysis, we employed 4 emitters instead of the minimum 2. The results are promising, revealing patterns analogous to those observed previously. For example, when plotting the cnot-vs-edge count diagram (Fig. S4), two distinct parts of the orbit become evident, with one showing a strong linear correlation. It is worth noting that the strength of the employed algorithm is in its effectiveness in depth reduction, thus the number of cnots, even in the best case (scaling as $3\times\lfloor n/3\rfloor$ with $n$ being the size of the graph), is larger compared to our previously identified optimal case ($n/2-2$). However, when considering circuit depth (to be specific emitter depth as defined in the main text) the results show an advantage over the circuit depths obtainable with the previous protocol ($n/2+6$ vs $n$ scaling). For the 4-emitter scenario, the optimal graph featuring the least depth coincides with the case that required the least number of cnots. Additionally, the results can be straightforwardly generalize to arbitrary large RGS as we did before; no matter the size of the RGS, using orbit enumeration Algorithm (1), the optimal shape resulting in the above mentioned advantages in circuit depth, is always the $3^{rd}$ graph in the list of graphs added to the orbit in the algorithm.
Figure S4: Number of cnots vs Number of edges (x-axis in logarithmic scale) correlation diagram of a 30-qubit RGS. Points correspond to graphs in the orbit. The orbit show two distinct behaviors, a group of members showing linear correlation while the other is constant in the number of cnots.
C: Repeater graph performance
A repeater graph is made up of a set of core nodes forming a complete sub-graph, and a set of leaf nodes (arms) attached to each one of the core qubits. The final goal of a repeater state can be thought of as distilling a perfect maximally entangled EPR pair that is established between two distance parties by distributing the leaf qubits (Fig. S5). Therefore, the overall performance of a single repeater graph state can be determined by evaluating the quality of this distilled 2-qubit state, that can be quantified, for instance, via its fidelity with a maximally entangled state.
Figure S5: EPR distillation process. From left to right, first a pair of qubits are selected to comprise the final 2-qubit state, the rest of the qubits, except the neighbors of the chosen pair are measured in the Z basis, and the two neighbors are measured in the X basis. The resultant state is locally equivalent to a graph where the two qubits are connected to each other, representing a maximally entangled state.
We now estimate this final fidelity for two cases, a 10-qubit and a 20-qubit RGS, and for each case we consider 2 scenarios: (i) using the original generation circuit and (ii) using the optimized one with reduced cnots. Considering the emitter-emitter CNOT gates between to be dominant sources of imperfection, we applied a noise model with an added depolarizing channel (with a depolarizing rate of 1%) to each of the CNOT gates in the generation circuit. For each case, we simulate the distillation procedure for different choices of the two arms used to form the EPR pair, and report the lower-bound for the fidelity of the final state. The results are listed in the table Table S1, and indicate that the same fidelity for the non-optimized scenario can be achieved by using the optimized circuit for an RGS twice the size.
Table S1: Fidelity of the final entangled pair distilled from a repeater graph state
           RGS size       non-optimized       optimized
       n=10       $88\pm 1\%$       $95\pm 1\%$
       n=20       $75\pm 1\%$       $88\pm 1\%$
Simulating for a range of sizes and depolarization rates, the results remain similar, confirming that the optimization allows for doubling the size of the generated RGS state without affecting the performance of the repeater. Such increase in size can lead to orders of magnitude increase in the achievable quantum key distribution rate. Specifically, the repeater protocol’s success rate is given by $(1-(1-{P_{b}}^{n/4}))^{N}$ [3]. Here, $n$ is number of photons in RGS, $N$ is the number of repeater stations between the two parties, and $P$ is the probability of a successful Bell state measurement between two photons from adjacent repeater stations. Using the parameters from [3]’s realistic example cases, and considering a total distance of $400$ $km$, with $N=100$ and $P_{b}=0.416$, we find an $2.66\times 10^{5}$ increase in success rate when $n$ is doubled from $15$ to $30$ qubits.
The value for $P_{b}$ is obtained by multiplying the probability of both photons making it to the Bell state measurement (BSM) station by the success probability of the BSM itself ($=50\%$). Assuming a loss of $\eta=0.2$ $dB/km$ and $2$ $km$ travel distance for each photon from two adjacent RGS to reach the BSM station in between (total distance $=400/100=4$ $km$) the probability of having both photons is equal to $0.832$ and then $P_{b}=0.416$.
D: Correlation analysis
In a correlation analysis of the local equivalency class of a graph, the goal is to determine whether there exists a significant correlation between a graph metric (a parameter related to the shape of a graph) and a cost parameter of its generating circuit. We examine how changes in the average values of one metric (chosen as the dependent variable) correspond to variations in the other metric (chosen as the independent variable) across the locally equivalent graphs in an orbit or in a sample from it. For instance, in Fig. S6, the independent variable (x-axis) the is set to be the number of edges in the graph and the dependent variable (y-axis) is chosen as the average of the circuit’s gate-count over the graphs sharing the same number of edges in the sample taken from the orbit. This diagram is derived for a sample of the LC orbit of a random graph of size $n$=20 with an edge probability of $p$=0.95. The Pearson correlation coefficient, used to measures the degree of linear relationship between two variables [4], for this specific diagram is $94\%$. Repeating the same procedure for 100 random graphs of the same type ($n$=20, $p$=0.95), we find the distribution of the correlation coefficients to be as seen in Fig. S7. This shows an average Pearson correlation of $0.86\pm 0.07$, indicating that for most of the graphs of this type, a strong correlation (over 0.5) holds true.
Figure S7: The distribution of Pearson coefficients for ‘edge count vs gate count’ correlations for 100 trials with different random graphs of size 20 and edge probability 0.95.
The edge reduction strategy used here take at most $\mathcal{O}(n^{2})$ steps, as in each step, we require for the number of edges to be reduced at least by one, and there are at most $n(n-1)/2$ edges in a graph. The clustering coefficient of the graph needs to be calculated in each step, which has a time complexity of $\mathcal{O}(n^{3})$ for a brute-force algorithm. Therefore, the total complexity will be upper bounded by $\mathcal{O}(n^{5})$. However, in practice, the actual scaling with size is much smaller than the worst case scenario, for example, in the specific use case of dense graphs (edge probability 95%), the scaling turns out to be $n^{2.68}$ for our python implementation (see Fig. S10). The method may thus be used to for optimization of graphs with thousands of qubits within minutes on commercial computers.
Figure S10: Polynomial scaling of the runtime with graph size for the proposed edge reduction strategy. Error bars are 1 standard deviation for data averaged over 100 trials.
E: Exhaustive analysis of small graphs
@@
To showcase the results of the exhaustive analysis, we look into an entanglement class associated with the graph (A) in Fig. S11, i.e., all graph states that are LC equivalent to this graph up to relabeling of the nodes. We have identified multiple intermediate target graphs to utilize in the generation process of any member of this entanglement class, each tailored for optimizing a certain cost parameter. In Fig. S11, examples of a case with the least number of cnots (B), a case with the smallest emitter depth (C), and a case with minimal number of unitary gates (D) are presented. More details for each option are listed in Table S2. We remark that there might be more than one best performing intermediate graphs depending on the cost function used. One can find the maximum improvement for each cost parameter by comparing the best values with the worst ones present in the last column, showing the maximum of the cost parameters over this class when no optimization is done. Overall, any arbitrary cost function can be employed to filter the results of such an exhaustive analysis, e.g, the fidelity of the final state with the target when a proper noise model is employed. The full dataset containing the results from the analysis of all 45 entanglement classes of graph with $n\leq 7$ nodes is available at Ref. [6].